Use the following information about Hospital X to answer the questions that follow (review chapter 5 for information about formulas used to calculate, and make sure to show all work for full credit) 

Fixed Cost = $5,000,000 

Variable Cost Per Patient Day = $500 

Revenue Per Patient Day = $750

2013 Patient Days =  34.000

 1. (1 pt) What are the total costs for the organization?

 2. (1 pt) Is the hospital making a profit or loss? What is the profit loss?

 3. (1pt) What is the total contribution margin at 34,000 patient days?
Answer #1

1) Calculation of total costs:

Total cost= fixed cost+ variable cost

Total cost= 5000000+(500*34000)

Total cost= $22000000

2) calculation of profit/loss:

Profit = revenue- total cost

Profit= (750*34000)-22000000

Profit= $3500000

3) Calculation of contribution:

Contribution= no. Of patients*(revenue-variable cost)

Contribution= 34000*(750-500)= $8500000